OpenStudy (anonymous):

I was stuck at number four where I have to multiply each terms by a constant that will make all coefficients integers. M y equation is 0=3.14x^3+6.28x^2-50.24 Height is 4inches and radius is 2inches. so I have (x+2) However, when I multiplied them all by 2, it still shows in decimals. IF I have multiplied them my 100, it looks like 0=314x^3+628x^2-5024. when I began to do (P/Q) P:5024 is pretty big number to find out the possibilities.. same as 314. Question: Apply the formula of a right circular cylinder to find the volume of the object. I have = radius=2inch, height=4inch.. volume=50.42 Basically my polynomial equation is 0=3.14x^3+6.28x^2-50.24 Since its decimals, I have to make it all coefficients integers.. it said in order to find them, multiply all by the constant, which is 2.....

OpenStudy (missmob):

..................................................

OpenStudy (missmob):

sharp wat

hartnn (hartnn):

5024/314 = 16

hartnn (hartnn):

so, possible roots will be factors of 16

hartnn (hartnn):

you could have divided everything by 3.14

OpenStudy (anonymous):

can't you just use the quadratic equation here? factoring will be annoying

OpenStudy (anonymous):

@hartnn will still be decimal

hartnn (hartnn):

x^3+2x^2-16=0

OpenStudy (anonymous):

\(6.26\ne2(3.14)\)

hartnn (hartnn):

aawww :(

hartnn (hartnn):

by any chace is it 6.28 ?

OpenStudy (anonymous):

more than likely meant to be :-)

OpenStudy (anonymous):

mayb I should write all the question so u may know what IM talking about. it s confusing right??

hartnn (hartnn):

yeah.

OpenStudy (anonymous):

ldrin.batak, u r right I better do quadratic eq. instead of factoring but I still have to find out the Fundamental Theorem of Algebra, the Rational Root theorem, Descartes`s rule of sign and the factor theorem......

hartnn (hartnn):

how you got this :my polynomial equation is 0=3.14x^3+6.26x^2-50.24 ? and 2pi means it has to be 6.28 ...and not 6.26...

OpenStudy (anonymous):

@LIZ-0- don't round to decimals yet... also are your powers of \(x\) correct?

OpenStudy (anonymous):

well i got this "my eq." by rewriting the equation with one variable woud result in a polynomial eq. that you could slove to find the radius... which it looks like... x=sqrt( (volume)/(piexheight))

OpenStudy (anonymous):

volume: v=sqrt( (v)/(piexheight)) so then when I rewrite the formula using the variable x for the radius.. which looks like.. x^2= ( (50.24)/ (pie) (x+2) ) subsitute..... pie(3.14)x^3+6.28x^2=50.24 this is wat I got.

OpenStudy (anonymous):

MY BIG MISTAKE it not 6.26 its 6.28 sry... I wrote 6.28 but I typed it in 6.26.... sry.......... hartnn and oldrin.bataku

hartnn (hartnn):

if h=x+2 \(V=\pi r^2 h \implies 50.24 =\pi r^2(r+2) \implies 50.24/\pi=r^3+2r^2\\ \implies r^3+2r^2-16=0\) try and solve this ...

hartnn (hartnn):

so, possible roots are factors of 16......try r=+2 or -2 ....

hartnn (hartnn):

oh, you already have it in the question :P r= 2inch, h= 2+2=4 inch

OpenStudy (anonymous):

sry Im really slow and don know wat 2 solve....

hartnn (hartnn):

what do you need to find out ?? the volume is already given to be 50.42 ....

OpenStudy (anonymous):

I have to find rational root theorem and decartes rule sign and the factor theorem.. now I am confused what is my polynomial eq. in order to find theses questions.......

hartnn (hartnn):

"find rational root theorem" ...how can one find a theorem ? :O ... you already found correct polynomial equation, x^3+2x^2-16=0

OpenStudy (anonymous):

It said Find the solution to this eq. algebraically using the Fundamental Theorem of Algebra, the Rational root theorem, decartes sign rule and factor theorem....

OpenStudy (anonymous):

fundamental theorem of algebra for this eq. x^3+2x^2-16=0,, there are 3 roots.degree... (is that a answer for this questions??)

OpenStudy (jack1):

The fundamental theorem of algebra states that every non-constant single-variable polynomial with complex coefficients has at least one complex root... This question... so many words...

OpenStudy (anonymous):

yup so many words, still it helps.

OpenStudy (jack1):

so your polynomial, one of the ways it can be factored is: x^3+2x^2-16=0 so = (x - 2)*(x^2 + 4x + 8) = 0 so one of the roots is x=2... as (2 - 2)*(x^2 + 4x + 8) = 0 ( 0 ) *(x^2 + 4x + 8) = 0 ( anything multiplied by zero = 0 ) so do you have to work out the other 2 roots?

OpenStudy (jack1):

cos they work out as complex numbers basically you're finding (x^2 + 4x + 8) = 0 so do you know how to find the x values that solve that...?

OpenStudy (anonymous):

wait Jack, but the constant is 2 and I have (x+2) but y (x-2) to conti. finding solutions??

OpenStudy (jack1):

...what? i didn't read everything above (as you oldrin and hartnan basically wrote a novel) y(x-2)... where did this come from?

OpenStudy (anonymous):

it come from x^3+2x^2-16=0 so = (x - 2)*(x^2 + 4x + 8) = 0 so one of the roots is x=2... as (2 - 2)*(x^2 + 4x + 8) = 0 ( 0 ) *(x^2 + 4x + 8) = 0 ( anything multiplied by zero = 0 ) it said (x-2) but yyyy?????

OpenStudy (jack1):

oh, ok, sorry i factorised x-2 out of your original polynomial x^3+2x^2-16 = (x - 2)*(x^2 + 4x + 8) same as basic factorizing of quadratics: 2x^2 - x -3 = (2x - 3) (x + 1)

OpenStudy (jack1):

and so to continue using this as an example: 2x^2 - x -3 = (2x - 3) (x + 1) so if 2x^2 - x -3 = 0 (2x - 3) (x + 1) = 0 so x = -1 or x = 1.5... yeah?

OpenStudy (jack1):

u ok with that liz?

OpenStudy (anonymous):

when I factorize, I have to do (x-2) instead of (x+2)

OpenStudy (jack1):

sorry is that a question or a statement?

OpenStudy (anonymous):

thats a question...sry

OpenStudy (jack1):

i suppose you could factorise x+2 out of it k so you got: ((x+2) * x^2) -16 = 0... that dosent factorise it enough to give you one of the roots u need 2 or 3 multiples, whereas you still have a "-16" in there by itself, so the order of operations wont give you 0 unless (x+2)* x^2 = 16... which isnt feasible or really easy to work out roots from that

OpenStudy (anonymous):

ok..

OpenStudy (jack1):

k so stick with x-2 as your first factor... so = 0 when x = +2 then do u need to work out the other 2 complex roots of your polynomial...? or move on to the other questions?

OpenStudy (anonymous):

oh yup... move on the other quest...

OpenStudy (jack1):

so it looks like you've already used descartes' to find the number of real roots... yeah?

OpenStudy (anonymous):

oh thats true..... !

OpenStudy (anonymous):

basically im finish???

OpenStudy (jack1):

so bam, one down rational root theorum... dude ive never seen this before, i can solve polynomials ok and vaguely understand them... but this is some serious maths ;)

OpenStudy (jack1):

yeah, if you dont need to find the complex roots of your poly

OpenStudy (jack1):

then ur pretty much done

OpenStudy (anonymous):

ok.. but wat abou fundamental algebra ... fundamental theorem of algebra...

OpenStudy (jack1):

that was the first one we were doing... find the roots of your poly... we found the real, the complex ones are found by continuing to solve: (x - 2)*(x^2 + 4x + 8) = 0 so we've got x-2 = 0 when x = +2 so all we need is the other half: (x^2 + 4x + 8) = 0 when... complex numbers are involved so complete the square then you end up with (x+2)^2 = -4 so roots are x+2 = 2i or x+2 = -2i so x = 2i -2 OR x = -2i -2

OpenStudy (jack1):

food, shower calling slaters liz

OpenStudy (anonymous):

sry Jack, but thnx!!Actually I understood wat u said!! thats really helps!! thnx! enjoy ur evening!